# Learning From Correctness Without Prompting Makes LLM Efficient Reasoner

### v1

@@ Abstract
Large language models (LLMs) have demonstrated outstanding performance across various tasks, yet they still exhibit limitations such as hallucination, unfaithful reasoning, and toxic content. One potential approach to mitigate these issues is learning from human or external feedback (e.g. tools). In this paper, we introduce an intrinsic self-correct reasoning framework for LLMs that eliminates the need for human feedback, external tools, and handcraft prompts. The proposed framework, based on a multi-step reasoning paradigm Learning from Correctness (LeCo), improves reasoning performance without needing to learn from errors. This paradigm prioritizes learning from correct reasoning steps, and a unique method to measure confidence for each reasoning step based on generation logits. Experimental results across various multi-step reasoning tasks demonstrate the effectiveness of the framework in improving reasoning performance with reduced token consumption.
@@ 1 Introduction
Large language models (LLMs) (OpenAI, 2023; Touvron et al., 2023) have exhibited remarkable performance on a diverse range of natural language processing benchmarks (Hendrycks et al., 2021a; Srivastava et al., 2022) and also showcased promising results on real-world applications (Wu et al., 2023; Thirunavukarasu et al., 2023). However, it is imperative to acknowledge that LLMs still possess certain limitations. For instance, the occurrence of undesirable behaviors like hallucinations (Rawte et al., 2023), generating harmful content (Bai et al., 2022), and non-adherence to established rules and constraints (Ouyang et al., 2022; Peng et al., 2023) remains largely unexplored.
One extensively employed approach to address these problems is learning from feedback (Pan et al., 2023). It involves guiding LLMs to improve their responses through a cycle of trial, examination, and correction. During the examination phase, feedback is provided to identify the shortcomings in the trial answer and guide the necessary corrections. Huang et al. (2023a); Gou et al. (2023a) have confirmed high-quality feedback can offer valuable insights into further corrections. Although human feedback (Ouyang et al., 2022; Fernandes et al., 2023) and external tools feedback (Gou et al., 2023a; b) are generally valuable, they are either expensive to collect or heavily dependent on the abilities of the selected tools. To eliminate external intervention, another popular line of research is self-correction, where the model progressively learns from the feedback it generates internally, without relying on external sources (An et al., 2023). However, Huang et al. (2023b) recently suggests that LLMs do not possess the inherent capabilities to find the errors and rectify their responses just by designing the prompts. More frustratingly, these methods often require creating extensive and elaborate handcraft prompts to guide the model in acquiring and understanding the feedback, which is a time-consuming and labor-intensive process, finally tuning our researchers into “prompt engineers”.
@@ 2 Related Work
Improving LLMs through learning from feedback has become a prevalent strategy, notably through reinforcement learning from human feedback, which seeks to align LLMs with human values by refining their outputs based on feedback (Ouyang et al., 2022; Bai et al., 2022; Touvron et al., 2023). However, this method faces challenges such as high costs due to manual labor and a lack of real-time feedback capabilities (Pan et al., 2023; Fernandes et al., 2023). An alternative strategy involves using self-correcting LLMs, which rely on automated feedback to iteratively adapt and understand the consequences of their actions without heavy reliance on human intervention. This feedback can be derived from outside sources such as other models (Yang et al., 2022; Lightman et al., 2023; Xiong et al., 2023), tools (Gou et al., 2023a; Charalambous et al., 2023), knowledge bases (Gao et al., 2023; Yu et al., 2023), or evaluation metrics (Jung et al., 2022; Welleck et al., 2023).
External feedback leverages external perspectives to identify errors and verify factual accuracy, offering insights that may not be recognized by the LLM alone. Conversely, feedback can also be internally generated, where the LLM evaluates and refines its output iteratively until a desired quality is achieved (Madaan et al., 2023; Shinn et al., 2023; Helbling et al., 2023; Xie et al., 2023). This self-improvement mechanism is particularly valuable in scenarios where external feedback is scarce or restricted (Yan et al., 2023; Ye et al., 2023). However, recent effort (Huang et al., 2023b) suggests that LLMs struggle to independently identify and correct errors through self-generated prompts. Unlike existing efforts, LeCo focuses on learning from one’s correct reasoning steps, without the need for feedback mechanisms including human intervention, external tools, or tailored prompts.
@@
Another strategy to facilitate reasoning involves instruction tuning, which leverages a significant volume of chain-of-thought (CoT) data (Chung et al., 2022; Mukherjee et al., 2023; Gunasekar et al., 2023; Luo et al., 2023). Recently, Liu et al. (2024) proposed to tune LLMs by comparing the logits differences between a pair of tuned and untuned smaller models, showcasing improvements in reasoning without CoT distillation. In contrast to these methods, our LeCo introduces an intrinsic self-correct reasoning mechanism that does not depend on fine-tuning or auxiliary models.
@@ 3.1 Step Confidence
Following the intra-step measurements, we sought to quantify the transition between consecutive steps. Our preliminary experiments yielded two key insights: 1) steps with lower overall confidence tend to have lower confidence levels specifically in the initial heading tokens (typically the first three). 2) These initial heading tokens were also the most likely to change across different program runs. Based on these observations, we propose using the probabilities of the heading tokens in a step to represent the inter-step transition score between that step and the subsequent one. In other words, the transition score is determined by:
@@
where $K$ is set to $3$ here.
@@ 5 Further Analyses
At the fine-grained level, we deeply investigate the design of step confidence, which involves calculating the sum of the average token confidence, step divergence score, and inter-step transition score. To minimize the time and token consumption, we employ the accuracy of identifying the earliest error step as our metric. This measurement has proven to be crucial for enhancing reasoning performance in subsequent rounds, as evidenced by the results in Table 5. To this end, we randomly sampled 100 incorrect solutions on the GSM8K dataset and manually annotated the earliest error step for these solutions. Then, we divide the predicted step into three categories, including exact_correct, partial_correct and wrong, wherein exact_correct means the predicted step is exactly the labeled earliest step; partial_correct means the predicted step is actually an error step but located after the earliest step, and wrong means the predicted step is before the target location. As presented in Table 5, LeCo performs best in finding the earliest error step, with accuracy over 50%. We also observe the significant performance drops when separately adopting one of these factors. More interestingly, among the three factors, we find the inter-step transition score affects the final performance most. This finding is also well-aligned with the observations in our preliminary experiments, as stated in Section 3.1, which suggests that the heading tokens of a step warrant more attention.
Figure 2: Evaluation of the changes after rethink stage. We compare our LeCO and RCI on GSM8K and StrategyQA datasets with GPT-3.5. W2R: the wrong answer is changed to right. R2W:the right answer is altered to wrong. W2W: a wrong answer is changed to another wrong answer. No change: The answer remains unchanged.
@@
As discussed above, the majority of initial solutions would not be modified after the rethink stage, which additionally escalates token consumption and ratio of “correct $\Rightarrow$ incorrect”. To alleviate these problems, we present an early stop strategy of LeCo, which dynamically determines whether the initial solution requires the refinement based on the overall solution score.
@@
Firstly, we conducted test on the GSM8K dataset using GPT-3.5-Turbo and recorded the solution confidence scores following Equation 7. As shown in Figure 3(a), we observed that the distributions of scores for both correct and incorrect solutions consistently tend to follow the norm distribution, with the average point of correct answers notably surpassing that of incorrect ones. We aim to employ this discrepancy to early stop the rethink stage. Specifically, we first randomly sample a subset from the testing data to obtain the distribution of solution scores, approximately 1/6 of the data of the entire test set used. Figure 3(b) illustrates the distribution on the GSM8K sample set, which also follows the norm distribution. Then, based on the 3-$\sigma$ characteristics of the norm distribution, we adopt the positive 1-$\sigma$ value from the score distribution of the incorrect solutions ($\mu+\sigma$) as our threshold, which covers 84% incorrect samples while only including around 50% correct instances.
As demonstrated in Table 7, consistent improvements can be obtained with early-stop LeCo over the vanilla CoT-based method. Compared to the standard LeCo, there are slight performance drops since more incorrect instances are filtered and not modified. However, early-stop LeCo can still maintain the performance levels intermediate to those of SC and LeCo while using fewer iteration rounds and tokens, approximately further reducing 10% tokens against the standard LeCo (More details in Appendix B). We note that early-stop LeCo is an alternative choice for the users to achieve a better trade-off between the token consumption and performance.
@@ A.2 Average Iterations Numbers by Different Methods and Models
Table 11 and 12 presents the average iteration numbers on the arithmetic reasoning, commonsense reasoning, logical reasoning and complex mathematical reasoning using OpenAI models. Table 13 illustrate the average iteration numbers on the GSM8K and MATH datasets using the DeepSeek model.
@@ B.1 Algorithm of Early stop LeCo
1:input questions $x$, model $M$, demonstration $Demo_{x}$, standard $*\textsc{LeCo}(\text{*})$, sample amount $R$, solution score $*sln\_score(\text{*})$, normailize function $*norm(\text{*})$
@@ B.2 Token Consumption and Iteration Number of Early Stop LeCo
Appendix C Case Study of LeCo
Table 16, 17 and 18 lists the specific cases of reasoning results via different methods on GSM8K, StrategyQA and MATH datasets.

### v2

@@ Abstract
Large language models (LLMs) have demonstrated outstanding performance across various tasks, yet they still exhibit limitations such as hallucination, unfaithful reasoning, and toxic content. One potential approach to mitigate these issues is learning from human or external feedback (e.g. tools). In this paper, we introduce an intrinsic self-correct reasoning framework for LLMs that eliminates the need for human feedback, external tools, and handcraft prompts. The proposed framework, based on a multi-step reasoning paradigm Learning from Correctness (LeCo), improves reasoning performance without needing to learn from errors. This paradigm prioritizes learning from correct reasoning steps, and a unique method to measure confidence for each reasoning step based on generation logits. Experimental results across various multi-step reasoning tasks demonstrate the effectiveness of the framework in improving reasoning performance with reduced token consumption. The code is available at https://github.com/starrYYxuan/LeCo.
@@ 1 Introduction
Large language models (LLMs; Brown et al. 2020; OpenAI 2023; Touvron et al. 2023) have exhibited remarkable performance on a diverse range of natural language processing benchmarks (Hendrycks et al., 2021a; Srivastava et al., 2022) and also showcased promising results on real-world applications (Wu et al., 2023; Thirunavukarasu et al., 2023). However, it is imperative to acknowledge that LLMs still possess certain limitations. For instance, the occurrence of undesirable behaviors like hallucinations (Rawte et al., 2023), generating harmful content (Bai et al., 2022), and non-adherence to established rules and constraints (Ouyang et al., 2022; Peng et al., 2023) remains largely unexplored.
One extensively employed approach to address these problems is learning from feedback (Pan et al., 2023). It involves guiding LLMs to improve their responses through a cycle of trial, examination, and correction. During the examination phase, feedback is provided to identify the shortcomings in the trial answer and guide the necessary corrections. Prior efforts (Huang et al., 2023a; Gou et al., 2023a) have confirmed high-quality feedback can offer valuable insights into further corrections. Although human feedback (Ouyang et al., 2022; Fernandes et al., 2023) and external tools feedback (Gou et al., 2023a; b) are generally valuable, they are either expensive to collect or heavily dependent on the abilities of the selected tools. To eliminate external intervention, another popular line of research is self-correction, where the model progressively learns from the feedback it generates internally, without relying on external sources (An et al., 2023). However, Huang et al. (2023b) recently suggests that LLMs do not possess the inherent capabilities to find the errors and rectify their responses just by designing the prompts. More frustratingly, these methods often require creating extensive and elaborate handcraft prompts to guide the model in acquiring and understanding the feedback, which is a time-consuming and labor-intensive process, finally tuning our researchers into “prompt engineers”.
@@ 2 Related Work
Improving LLMs through learning from feedback has become a prevalent strategy, notably through reinforcement learning from human feedback, which seeks to align LLMs with human values by refining their outputs based on feedback (Ouyang et al., 2022; Bai et al., 2022; Touvron et al., 2023). However, this method faces challenges such as high costs due to manual labor and a lack of real-time feedback capabilities (Pan et al., 2023; Fernandes et al., 2023). An alternative strategy involves using self-correcting LLMs, which rely on automated feedback to iteratively adapt and understand the consequences of their actions without heavy reliance on human intervention. This feedback can be derived from outside sources such as other models (Yang et al., 2022; Lightman et al., 2023; Xiong et al., 2023), tools (Huang et al., 2024; Lu et al., 2024b), knowledge bases (Gao et al., 2023; Yu et al., 2023), or evaluation metrics (Jung et al., 2022; Welleck et al., 2023).
External feedback leverages external perspectives to identify errors and verify factual accuracy, offering insights that may not be recognized by the LLM alone. Conversely, feedback can also be internally generated, where the LLM evaluates and refines its output iteratively until a desired quality is achieved (Madaan et al., 2023; Shinn et al., 2023; Helbling et al., 2023; Xie et al., 2023). This self-improvement mechanism is particularly valuable in scenarios where external feedback is scarce or restricted (Yan et al., 2023; Lu et al., 2024a). However, Huang et al. (2023b) suggests that LLMs struggle to independently identify and correct errors through self-generated prompts. Recent effort (Gonen et al., 2023) show that an LLM’s familiarity with a prompt’s language predicts its effectiveness, with lower perplexity prompts leading to better performance. Unlike existing efforts, LeCo focuses on learning from one’s correct reasoning steps, without the need for feedback mechanisms including human intervention, external tools, or tailored prompts.
@@
Another strategy to facilitate reasoning involves instruction tuning, which leverages a significant volume of chain-of-thought (CoT) data (Chung et al., 2022; Mukherjee et al., 2023; Gunasekar et al., 2023; Luo et al., 2023). Recently, Liu et al. (2024) proposed to tune LLMs by comparing the logit differences between a pair of tuned and untuned smaller models, showcasing improvements in reasoning without CoT distillation. In contrast to these methods, our LeCo introduces an intrinsic self-correct reasoning mechanism that does not depend on fine-tuning or auxiliary models.
@@ 3.1 Step Confidence
Following the intra-step measurements, we sought to quantify the transition between consecutive steps. Our preliminary experiments yielded two key insights: 1) steps with lower overall confidence tend to have lower confidence levels specifically in the initial heading tokens (typically the first three), more dicussions can be found at Section D. 2) These initial heading tokens were also the most likely to change across different program runs. Based on these observations, we propose using the probabilities of the heading tokens in a step to represent the inter-step transition score between that step and the subsequent one. In other words, the transition score is determined by:
@@
where $K$ is set to $3$ here. Further analysis of hyperparameter settings are discussed in Section C
@@ 5 Further Analyses
At the fine-grained level, we deeply investigate the design of step confidence, which involves calculating the sum of the average token confidence, step divergence score, and inter-step transition score. To minimize the time and token consumption, we employ the accuracy of identifying the earliest error step as our metric. This measurement has proven to be crucial for enhancing reasoning performance in subsequent rounds, as evidenced by the results in Table 5. To this end, we randomly sampled 100 incorrect solutions on the GSM8K dataset and manually annotated the earliest error step for these solutions. Then, we divide the predicted step into three categories, including exact_correct, partial_correct and wrong, wherein exact_correct means the predicted step is exactly the labeled earliest step; partial_correct means the predicted step is an error step but located after the earliest step, and wrong means the predicted step is before the target location. As presented in Table 5, LeCo performs best in finding the earliest error step, with accuracy over 50%. We also observe the significant performance drops when separately adopting one of these factors. More interestingly, among the three factors, we find the inter-step transition score affects the final performance most. This finding is also well-aligned with the observations in our preliminary experiments, as stated in Section 3.1, which suggests that the heading tokens of a step warrant more attention.
Figure 2: Evaluation of the changes after the rethink stage. We compare our LeCO and RCI on GSM8K and StrategyQA datasets with GPT-3.5. W2R: the wrong answer is changed to right. R2W: the right answer is altered to wrong. W2W: a wrong answer is changed to another wrong answer. No change: The answer remains unchanged.
@@
As discussed above, the majority of initial solutions would not be modified after the rethink stage, which additionally escalates token consumption and ratio of “correct $\Rightarrow$ incorrect”. To alleviate these problems, we present an early stop strategy of LeCo, which dynamically determines whether the initial solution requires refinement based on the overall solution score.
@@
Firstly, we conducted the test on the GSM8K dataset using GPT-3.5-Turbo and recorded the solution confidence scores following Equation 7. As shown in Figure 3(a), we observed that the distributions of scores for both correct and incorrect solutions consistently tend to follow the norm distribution, with the average point of correct answers notably surpassing that of incorrect ones. We aim to employ this discrepancy to early stop the rethink stage. Specifically, we first randomly sample a subset from the testing data to obtain the distribution of solution scores, approximately 1/6 of the data of the entire test set used. Figure 3(b) illustrates the distribution on the GSM8K sample set, which also follows the norm distribution. Then, based on the 3-$\sigma$ characteristics of the norm distribution, we adopt the positive 1-$\sigma$ value from the score distribution of the incorrect solutions ($\mu+\sigma$) as our threshold, which covers 84% incorrect samples while only including around 50% correct instances.
As demonstrated in Table 7, consistent improvements can be obtained with early-stop LeCo over the vanilla CoT-based method. Compared to the standard LeCo, there are slight performance drops since more incorrect instances are filtered and not modified. However, early-stop LeCo can still maintain the performance levels intermediate to those of SC and LeCo while using fewer iteration rounds and tokens, approximately further reducing 10% tokens against the standard LeCo (More details in Appendix B). We note that early-stop LeCo is an alternative choice for the users to achieve a better trade-off between token consumption and performance.
@@ A.2 Average Iterations Numbers by Different Methods and Models
Table 11 and 12 present the average iteration numbers on arithmetic reasoning, commonsense reasoning, logical reasoning, and complex mathematical reasoning using OpenAI models. Table 13 illustrates the average iteration numbers on the GSM8K and MATH datasets using the DeepSeek model.
@@ B.1 Algorithm of Early stop LeCo
1:input questions $x$, model $M$, demonstration $Demo_{x}$, standard $*\textsc{LeCo}(\text{*})$, sample amount $R$, solution score $*sln\_score(\text{*})$, normalize function $*norm(\text{*})$
@@ B.2 Token Consumption and Iteration Number of Early Stop LeCo
Appendix C Hyperparameter Settings
We compared the experimental results under different settings and found that our method is relatively insensitive to hyperparameters, such as $K$ and $\tau$. We attach the experimental results of GPT-3.5 on GSM8K as follows.
K 1 3 5
Complex 81.8 80.89 83
LeCo + Complex 82.83 82.33 83.87
(+1.03) (+1.44) (+0.87)
Table 16: Settings of Hyperparameter $K$
$\tau$ 0.1 0.2 0.3 0.4 0.5
Complex 81.16 80.98 80.89 82.46 83.03
LeCo+Complex 82.46 82.24 82.33 83.88 83.84
(+1.3) (+1.26) (+1.44) (+1.42) (+0.81)
Table 17: Settings of Hyperparameter $\tau$
Table 16 and Table 17 present the settings of hyperparameter $K$ and $\tau$.
In the design of the transition score, the parameter $K$ determines the usage of several initial tokens, hence the value of $K$ can not be very large and we set $K$ varying from 1 to 5.
In the design of the divergence score, the parameter $\tau$ is used to rescale the KL divergence to a reasonable range and helps the divergence score to show significant performance. When $\tau$ exceeds 0.5 in the logarithmic function, the divergence diminishes to negligible values, such as 0.002 or 0.004, which fail to capture the desired differences. Consequently, our study focuses on the impact of $\tau$ within the range of 0.1 to 0.5.
The results, as depicted in the tables, reveal a consistent improvement, indicating the robustness of our method to these parameter.
Appendix D Preliminary Experiments
We draw the scatter plot of the relationship between the overall confidence score and inter-step transition score for 1000 reasoning steps. As shown in Fig4, it’s obvious that the overall confidence and inter-step transition scores are highly positively correlated.
Figure 4: The relation between overall confidence and inter-step transition scores
Appendix E Case Study of LeCo
Table 18, 19 and 20 lists the specific cases of reasoning results via different methods on GSM8K, StrategyQA, and MATH datasets.